OpenStudy (anonymous):

If the circumference of a circle is 12, and the angle measure of an arc is 60o, which is the length of the arc?

OpenStudy (jh3power):

12 times 60/360=arc so it's 12 times 1/6=2 it's2

OpenStudy (anonymous):

thanks

OpenStudy (jh3power):

welcome

OpenStudy (jh3power):

IT's generally 2pi r times theta/360=arc

OpenStudy (mathstudent55):

\(60^o\) is \( \frac{1}{6} \) of a circle \( (360^o)\), so you need \( \frac{1}{6}\) of 12.